Consider for example a hole being created in the Hyperloop tunnel, which leads to a sudden increase in the local air pressure from 99~\mbox{Pa}99 Pa to 101,325~\mbox{Pa}101,325 Pa while maintaining constant temperature and volume. If the normal drag force on the Hyperloop is 320~\mbox{N}320 N, how much acceleration in g's would the passengers in the Hyperloop experience if the pod hit the region of high pressure?

The mass of the pod is approximately 15000~\mbox{kg}15000 kg.
g=9.8~\mbox{m/s}^2g=9.8 m/s
2
1
Expert's answer
2020-04-23T13:02:35-0400

As per the question,

Initial pressure =99Pa

Final pressure =101,325 Pa

drag force = 320N

mass = 15000kg

Force unit area "F=(P_{final}-P_{initial })=101325-99=101226N"

Let the acceleration, a,

So, "15000g-101226-325N=15000a"

"a=3.03 m\/sec^2"